OpenStudy (atlchic):

solve by elimination, 2x-y=3 and 2/3x-y=-1.

OpenStudy (atlchic):

do elimination, plz

OpenStudy (anonymous):

2x - y = 3 -2/3x + y = +1 --------------- 4/3x = 4 x = 3 2x - y = 3 2*3 - y = 3 2*3 - 3 = y 3 = y